OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (f + g)(-1)?

OpenStudy (anonymous):

7

OpenStudy (saifoo.khan):

7 is 100% correct answer!

OpenStudy (anonymous):

haha i feel like youre being sarcastic but how did you get that number

OpenStudy (saifoo.khan):

-2x + 8 + 3x x + 8 (-1) + 8 7

OpenStudy (anonymous):

f+g=-2x + 8 + 3x =x + 8 (f+g)(-1)=7